# On the uniqueness of compiling graphs under the parity transformation

### v2

@@ 4.2 Uniqueness of compiled hypergraphs in one or two dimensions
$V_{e_{n},\ell}\coloneqq\left\{v\in V_{G_{H_{n-1},\ell_{n-1}}}\left|\,\exists e^{\prime}\in E_{n-1}\colon\exists v_{0}\in e^{\prime}\colon\exists v_{1}\in\tilde{e}_{n}\colon v\in\ell(v_{0})\cap\ell(v_{1})\right.\right\},$ (10)
@@
$\ell_{1}(e_{2}\setminus e_{1})$ =$=\left\{\left\{i_{(k_{1}+p-1\,\mathrm{mod}\,m)+1},i_{1}^{\prime}\right\},\left\{i_{1}^{\prime},i_{2}^{\prime}\right\},\ldots,\left\{i_{\left|e_{2}\right|-p-1}^{\prime},i_{k_{1}}\right\}\right\},$
$\ell_{1}(e_{1}\cap e_{2})$ =$=\left\{\left\{i_{k_{1}},i_{(k_{1}\,\mathrm{mod}\,m)+1}\right\},\ldots,\left\{i_{(k_{1}+p-2\,\mathrm{mod}\,m)+1},i_{(k_{1}+p-1\,\mathrm{mod}\,m)+1}\right\}\right\},$
@@
$\ell_{2}(e_{2}\setminus e_{1})$ =$=\left\{\left\{j_{(k_{2}+p-1\,\mathrm{mod}\,m)+1},j_{1}^{\prime}\right\},\left\{j_{1}^{\prime},j_{2}^{\prime}\right\},\ldots,\left\{j_{\left|e_{2}\right|-p-1}^{\prime},j_{k_{2}}\right\}\right\},,$
$\ell_{2}(e_{1}\cap e_{2})$ =$=\left\{\left\{j_{k_{2}},j_{(k_{2}\,\mathrm{mod}\,m)+1}\right\},\ldots,\left\{j_{(k_{2}+p-2\,\mathrm{mod}\,m)+1},j_{(k_{2}+p-1\,\mathrm{mod}\,m)+1}\right\}\right\}.$
@@ 4.3 Counterexamples for higher dimensions
Let $H=(V,E)$ be the hypergraph wit vertex set $V\coloneqq\left\{1,2,\ldots,10\right\}$ and edge set $E\coloneqq\left\{\left\{1,2,3,4\right\},\left\{2,5,6,7\right\},\left\{6,8,9,10\right\}\right\}$. Then, we choose the two loop labelings $\ell_{1}$ and $\ell_{2}$ according to Figure 9 and 10. Since $\dim(\mathcal{L}_{G_{H,\ell_{1}}})=\dim(\mathcal{L}_{G_{H,\ell_{2}}})=3$ we deduce from Theorem 3.3 that $[G_{H,\ell_{1}}],[G_{H,\ell_{2}}]\in\mathrm{par}|_{\mathscr{D}}^{-1}(\left\{y\right\})$ where $y\coloneqq\mathrm{par}([G_{H,\ell_{1}}])$. However, we observe that the degree of vertex $3$ in $H_{1}$ equals four whereas every vertex in $H_{2}$ has degree three at most. This implies that $G_{H_{1},\ell_{1}}$ and $G_{H_{2},\ell_{2}}$ are not isomorphic graphs. Figure 11 and 12 show the induced graphs of the loop labelings $\ell_{1}$ and $\ell_{2}$. We emphasize that the first loop labeling $\ell_{1}$ is constructed in such a way that vertex $3$ in the induced graph $G_{H,\ell_{1}}$ is contained in the labelings of the two vertices $2$ and $6$ in the four-body plaquette layout which connect all three plaquettes. Note that by iteratively appending new four-body plaquettes to the rightmost plaquette and extending the two labelings $\ell_{1}$ and $\ell_{2}$, counterexamples with an arbitrary number of plaquettes $n\geq 3$ can be constructed.
@@
We consider the hypergraph $H$ as illustrated in Figures 13 and 14 with seven edges and $16$ vertices and define the two loop labelings $\ell_{1}$ and $\ell_{2}$ as also presented in Figures 13 and 14. From Theorem 2.9 we deduce that $\dim(\mathcal{L}_{G_{H,\ell_{1}}})=\dim(\mathcal{L}_{G_{H,\ell_{2}}})=7$ and therefore by Theorem 3.3 we can conclude $[G_{H,\ell_{1}}],[G_{H,\ell_{2}}]\in\mathrm{par}|_{\mathscr{D}}^{-1}(\left\{y\right\})$ with $y\coloneqq\mathrm{par}([G_{H,\ell_{1}}])$. Similarly as in the previous example, we observe that the degree of vertex $3$ in $H_{2}$ equals six whereas every vertex in $H_{1}$ has degree five at most. This implies again that $G_{H_{1},\ell_{1}}$ and $G_{H_{2},\ell_{2}}$ are not isomorphic graphs. We note that in the edge $\left\{5,8,9,10\right\}$ we switched the labelings of the two vertices $9$ and $10$ in order to obtain a higher degree of the vertex $3$ in $G_{H,\ell_{2}}$.
@@ 4.4.1 Characterization of preimages of rectangular plaquette layouts
the “vertical” enumeration of $E$, where $k=1,\ldots,(m-1)(n-1)$. Here, ///// / denotes the floor division of integers. Then, we observe for $i\in\left\{1,2\right\}$
@@
    Now, we fix the elements $i_{1},\ldots,i_{4}\in\mathbb{N}$ in (i) and show by induction that there exist for all $12\leq k\leq n-$ 1 pairwise distinct numbers $i_{5},\ldots,i_{k+3}\in\mathbb{N}\setminus\left\{i_{1},i_{2},i_{3},i_{4}\right\}$ with
@@
    For $k=2$, applying Lemma 4.6 (i) with $H_{2,1}$ shows that there exists $i_{5}\in\mathbb{N}\setminus\left\{i_{1},i_{2},i_{3},i_{4}\right\}$ such that ($\ell(\left\{v_{1,3},v_{2,3}\right\})=\left\{\left\{i_{2},i_{5}\right\},\left\{i_{5},i_{4}\right\}\right\}$. Since $e_{m}^{(2)}\cap e_{m+1}^{(2)}=e_{1,2}\cap e_{2,2}=\left\{v_{2,2},v_{2,3}\right\}$, we obtain by (i) as well as applying Lemma 4.6 (ii) on $H_{m+1,2}$ the relations $\ell(v_{1,3})=\left\{i_{2},i_{5}\right\}$ and $\ell(v_{2,3})=\left\{i_{4},i_{5}\right\}$. The remaining statement follows by induction and the same arguments.
@@
    In the last step of the proof, we apply (i) and (ii) and choose pairwise distinct $i_{1},\ldots,i_{n+2}\in\mathbb{N}$ satisfying the first two conditions in Lemma 4.12. We finally show by induction that for all $12\leq k\leq m-$ 1 there exist pairwise distinct $i_{n+3},\ldots,i_{n+k+1}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+2}\right\}$ satisfying
@@
    Suppose that $k=2$. We use the same arguments as before and apply Lemma 4.6 (i) on $H_{2,2}$ to find $i_{n+3}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+2}\right\}$ with ($\ell(\left\{v_{3,1},v_{3,2}\right\})=\left\{\left\{i_{1},i_{n+3}\right\},\left\{i_{3},i_{n+3}\right\}\right\}$. Moreover, applying Lemma 4.6 (ii) on $H_{n+1,1}$ and using the first two conditions in Lemma 4.12, we deduce that $\ell(v_{3,2})=\left\{i_{3},i_{n+3}\right\}$ and therefore $\ell(v_{3,1})=\left\{i_{1},i_{n+3}\right\}$, which shows that (14) holds for $k=2$. Since $e_{2,2}=\left\{v_{2,2},v_{2,3},v_{3,3},v_{3,2}\right\}$ we deduce from $\ell(e_{2,2})\in L_{G_{H,\ell}}$ and the relations $\ell(v_{2,2})=\left\{i_{3},i_{4}\right\}$, $\ell(v_{3,2})=\left\{i_{3},i_{n+3}\right\}$, $\ell(v_{2,3})=\left\{i_{4},i_{5}\right\}$ that $\ell(v_{3,3})=\left\{i_{5},i_{n+3}\right\}$. This shows that (15) holds for $j=2$. Now, suppose $n\geq 4$ and (15) holds for some particular $12\leq j<n-$ 1 and $i=2$. By using the induction hypothesis, we observe that $\ell(v_{3,j+1})=\left\{i_{n+3},i_{j+3}\right\}$. Since $\ell(v_{2,j+1})=\left\{i_{4},i_{j+3}\right\}$ and $\ell(v_{2,j+2})=\left\{i_{4},i_{j+4}\right\}$ by (13), and $\ell(e_{2,j+1})\in L_{G_{H,\ell}}$ we therefore obtain $\ell(v_{3,j+2})=\left\{i_{n+3},i_{j+4}\right\}$.
    Assume now that $m\geq 4$ and the above statement with conditions (14) and (15) holds for a particular $12\leq k<m-$ 1, and let $i_{n+3},\ldots,i_{n+k+1}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+2}\right\}$ be some integers satisfying these two conditions. In order to show the statement for $k+1$, we are only left to show that there exists $i_{n+k+2}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+k+1}\right\}$ which fulfills conditions (14) and (15) for $i=k+1$. This again follows by induction and the same arguments.∎
@@
      ($\ell(\left\{v_{1,1},v_{2,1}\right\})=\left\{\left\{i_{1},i_{2}\right\},\left\{i_{1},i_{4}\right\}\right\}$,
@@
      ($\ell(\left\{v_{1,2},v_{2,2}\right\})=\left\{\left\{i_{2},i_{3}\right\},\left\{i_{3},i_{4}\right\}\right\}$ and
@@
    For $k=3$, we observe from the application of Lemma 4.6 (ii) on $H_{2}=(V_{2},E_{2})$ with $E_{2}=\left\{e_{1,1},e_{1,2}\right\}$ and $V_{2}=\left\{v\in V\mid v\in E_{2}\right\}$ that there exist pairwise distinct $i_{2},i_{3},i_{4}\in\mathbb{N}$ such that ($\ell(\left\{v_{1,2},v_{2,2}\right\})=\left\{\left\{i_{2},i_{3}\right\},\left\{i_{3},i_{4}\right\}\right\}$. Moreover, since $\ell(e_{1,1}),\ell(e_{1,2})\in L_{G_{H,\ell}}$ we find by Lemma 3.10 (ii) distinct numbers $i_{1},i_{5}\in\mathbb{N}\setminus\left\{i_{2},i_{3},i_{4}\right\}$ satisfying ($\ell(\left\{v_{1,1},v_{2,1}\right\})=\left\{\left\{i_{1},i_{2}\right\},\left\{i_{1},i_{4}\right\}\right\}$ and ($\ell(\left\{v_{1,3},v_{2,3}\right\})=\left\{\left\{i_{2},i_{5}\right\},\left\{i_{4},i_{5}\right\}\right\}$. Therefore, conditions (i)(a), (i)(b) and (i)(c) hold for $k=3$. Now, suppose $n\geq 4$ and the statement holds for some particular $3\leq k<n$. Since ($\ell(\left\{v_{1,k},v_{2,k}\right\})=\left\{\left\{i_{2},i_{k+2}\right\},\left\{i_{k+2},i_{4}\right\}\right\}$ we can find by Lemma 4.6 (i) a positive integer $i_{k+3}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{k+2}\right\}$ satisfying the relation (${\ell(\left\{v_{1,{k+1}},v_{2,{k+1}}\right\})=\left\{\left\{i_{2},i_{k+3}\right\},\left\{i_{k+3},i_{4}\right\}\right\}}$. Thus, by our induction hypothesis (i)(a), (i)(b) and (i)(c) hold for $k+1$, which shows the desired statement. As a consequence, we can find for the loop labelings $\ell_{1}$ and $\ell_{2}$ pairwise distinct numbers $i_{1},\ldots,i_{n+2}\in\mathbb{N}$ and $j_{1},\ldots,j_{n+2}\in\mathbb{N}$, respectively, satisfying conditions (i)(a)(i)(c) for $k=n$. Thus, we see that $V_{G_{H,\ell_{1}}}=\left\{i_{1},\ldots,i_{n+2}\right\}$ as well as $V_{G_{H,\ell_{2}}}=\left\{j_{1},\ldots,j_{n+2}\right\}$. Moreover, from the conditions (i)(a)(i)(c) we immediately obtain that the function $f\colon V_{G_{H,\ell_{1}}}\to V_{G_{H,\ell_{2}}}$ defined by $f(i_{k})\coloneqq j_{k}$ for $k=1,\ldots,n+2$ fulfills property (2), which shows that the induced graphs $G_{H,\ell_{1}}$ and $G_{H,\ell_{2}}$ are isomorphic.

### v3

@@ 4.2 Uniqueness of compiled hypergraphs in one or two dimensions
[…]
@@
$\ell_{1}(e_{2}\setminus e_{1})$ $=\left\{\left\{i_{(k_{1}+p-1\,\mathrm{mod}\,m)+1},i_{1}^{\prime}\right\},\left\{i_{1}^{\prime},i_{2}^{\prime}\right\},\ldots,\left\{i_{\left|e_{2}\right|-p-1}^{\prime},i_{k_{1}}\right\}\right\},$
$\ell_{1}(e_{1}\cap e_{2})$ $=\left\{\left\{i_{k_{1}},i_{(k_{1}\,\mathrm{mod}\,m)+1}\right\},\ldots,\left\{i_{(k_{1}+p-2\,\mathrm{mod}\,m)+1},i_{(k_{1}+p-1\,\mathrm{mod}\,m)+1}\right\}\right\},$
@@
$\ell_{2}(e_{2}\setminus e_{1})$ $=\left\{\left\{j_{(k_{2}+p-1\,\mathrm{mod}\,m)+1},j_{1}^{\prime}\right\},\left\{j_{1}^{\prime},j_{2}^{\prime}\right\},\ldots,\left\{j_{\left|e_{2}\right|-p-1}^{\prime},j_{k_{2}}\right\}\right\},,$
$\ell_{2}(e_{1}\cap e_{2})$ $=\left\{\left\{j_{k_{2}},j_{(k_{2}\,\mathrm{mod}\,m)+1}\right\},\ldots,\left\{j_{(k_{2}+p-2\,\mathrm{mod}\,m)+1},j_{(k_{2}+p-1\,\mathrm{mod}\,m)+1}\right\}\right\}.$
@@ 4.3 Counterexamples for higher dimensions
Let $H=(V,E)$ be the hypergraph with vertex set $V\coloneqq\left\{1,2,\ldots,10\right\}$ and edge set $E\coloneqq\left\{\left\{1,2,3,4\right\},\left\{2,5,6,7\right\},\left\{6,8,9,10\right\}\right\}$. Then, we choose the two loop labelings $\ell_{1}$ and $\ell_{2}$ according to Figure 9 and 10. Since $\dim(\mathcal{L}_{G_{H,\ell_{1}}})=\dim(\mathcal{L}_{G_{H,\ell_{2}}})=3$ we deduce from Theorem 3.3 that $G_{H,\ell_{1}}],[G_{H,\ell_{2}}]\in\mathrm{par}|_{\mathscr{D}}^{-1}(\left\{y\right\})$ where $y\coloneqq\mathrm{par}([G_{H,\ell_{1}}])$. However, we observe that the degree of vertex $3$ in $H_{1}$ equals four whereas every vertex in $H_{2}$ has degree three at most. This implies that $G_{H_{1},\ell_{1}}$ and $G_{H_{2},\ell_{2}}$ are not isomorphic graphs. Figure 11 and 12 show the induced graphs of the loop labelings $\ell_{1}$ and $\ell_{2}$. We emphasize that the first loop labeling $\ell_{1}$ is constructed in such a way that vertex $3$ in the induced graph $G_{H,\ell_{1}}$ is contained in the labelings of the two vertices $2$ and $6$ in the four-body plaquette layout which connect all three plaquettes. Note that by iteratively appending new four-body plaquettes to the rightmost plaquette and extending the two labelings $\ell_{1}$ and $\ell_{2}$, counterexamples with an arbitrary number of plaquettes $n\geq 3$ can be constructed.
@@
We consider the hypergraph $H$ as illustrated in Figures 13 and 14 with seven edges and $16$ vertices and define the two loop labelings $\ell_{1}$ and $\ell_{2}$ as also presented in Figures 13 and 14. From Theorem 2.9 we deduce that $\dim(\mathcal{L}_{G_{H,\ell_{1}}})=\dim(\mathcal{L}_{G_{H,\ell_{2}}})=7$ and therefore by Theorem 3.3 we can conclude $G_{H,\ell_{1}}],[G_{H,\ell_{2}}]\in\mathrm{par}|_{\mathscr{D}}^{-1}(\left\{y\right\})$ with $y\coloneqq\mathrm{par}([G_{H,\ell_{1}}])$. Similarly as in the previous example, we observe that the degree of vertex $3$ in $H_{2}$ equals six whereas every vertex in $H_{1}$ has degree five at most. This implies again that $G_{H_{1},\ell_{1}}$ and $G_{H_{2},\ell_{2}}$ are not isomorphic graphs. We note that in the edge $\left\{5,8,9,10\right\}$ we switched the labelings of the two vertices $9$ and $10$ in order to obtain a higher degree of the vertex $3$ in $G_{H,\ell_{2}}$.
@@ 4.4.1 Characterization of preimages of rectangular plaquette layouts
the “vertical” enumeration of $E$, where $k=1,\ldots,(m-1)(n-1)$. Here, //// denotes the floor division of integers. Then, we observe for $i\in\left\{1,2\right\}$
@@
    Now, we fix the elements $i_{1},\ldots,i_{4}\in\mathbb{N}$ in (i) and show by induction that there exist for all $2\leq k\leq n-1$ pairwise distinct numbers $i_{5},\ldots,i_{k+3}\in\mathbb{N}\setminus\left\{i_{1},i_{2},i_{3},i_{4}\right\}$ with
@@
    For $k=2$, applying Lemma 4.6 (i) with $H_{2,1}$ shows that there exists $i_{5}\in\mathbb{N}\setminus\left\{i_{1},i_{2},i_{3},i_{4}\right\}$ such that $\ell(\left\{v_{1,3},v_{2,3}\right\})=\left\{\left\{i_{2},i_{5}\right\},\left\{i_{5},i_{4}\right\}\right\}$. Since $e_{m}^{(2)}\cap e_{m+1}^{(2)}=e_{1,2}\cap e_{2,2}=\left\{v_{2,2},v_{2,3}\right\}$, we obtain by (i) as well as applying Lemma 4.6 (ii) on $H_{m+1,2}$ the relations $\ell(v_{1,3})=\left\{i_{2},i_{5}\right\}$ and $\ell(v_{2,3})=\left\{i_{4},i_{5}\right\}$. The remaining statement follows by induction and the same arguments.
@@
    In the last step of the proof, we apply (i) and (ii) and choose pairwise distinct $i_{1},\ldots,i_{n+2}\in\mathbb{N}$ satisfying the first two conditions in Lemma 4.12. We finally show by induction that for all $2\leq k\leq m-1$ there exist pairwise distinct $i_{n+3},\ldots,i_{n+k+1}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+2}\right\}$ satisfying
@@
    Suppose that $k=2$. We use the same arguments as before and apply Lemma 4.6 (i) on $H_{2,2}$ to find $i_{n+3}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+2}\right\}$ with $\ell(\left\{v_{3,1},v_{3,2}\right\})=\left\{\left\{i_{1},i_{n+3}\right\},\left\{i_{3},i_{n+3}\right\}\right\}$. Moreover, applying Lemma 4.6 (ii) on $H_{n+1,1}$ and using the first two conditions in Lemma 4.12, we deduce that $\ell(v_{3,2})=\left\{i_{3},i_{n+3}\right\}$ and therefore $\ell(v_{3,1})=\left\{i_{1},i_{n+3}\right\}$, which shows that (14) holds for $k=2$. Since $e_{2,2}=\left\{v_{2,2},v_{2,3},v_{3,3},v_{3,2}\right\}$ we deduce from $\ell(e_{2,2})\in L_{G_{H,\ell}}$ and the relations $\ell(v_{2,2})=\left\{i_{3},i_{4}\right\}$, $\ell(v_{3,2})=\left\{i_{3},i_{n+3}\right\}$, $\ell(v_{2,3})=\left\{i_{4},i_{5}\right\}$ that $\ell(v_{3,3})=\left\{i_{5},i_{n+3}\right\}$. This shows that (15) holds for $j=2$. Now, suppose $n\geq 4$ and (15) holds for some particular $2\leq j<n-1$ and $i=2$. By using the induction hypothesis, we observe that $\ell(v_{3,j+1})=\left\{i_{n+3},i_{j+3}\right\}$. Since $\ell(v_{2,j+1})=\left\{i_{4},i_{j+3}\right\}$ and $\ell(v_{2,j+2})=\left\{i_{4},i_{j+4}\right\}$ by (13), and $\ell(e_{2,j+1})\in L_{G_{H,\ell}}$ we therefore obtain $\ell(v_{3,j+2})=\left\{i_{n+3},i_{j+4}\right\}$.
    Assume now that $m\geq 4$ and the above statement with conditions (14) and (15) holds for a particular $2\leq k<m-1$, and let $i_{n+3},\ldots,i_{n+k+1}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+2}\right\}$ be some integers satisfying these two conditions. In order to show the statement for $k+1$, we are only left to show that there exists $i_{n+k+2}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{n+k+1}\right\}$ which fulfills conditions (14) and (15) for $i=k+1$. This again follows by induction and the same arguments.∎
@@
      $\ell(\left\{v_{1,1},v_{2,1}\right\})=\left\{\left\{i_{1},i_{2}\right\},\left\{i_{1},i_{4}\right\}\right\}$,
@@
      $\ell(\left\{v_{1,2},v_{2,2}\right\})=\left\{\left\{i_{2},i_{3}\right\},\left\{i_{3},i_{4}\right\}\right\}$ and
@@
    For $k=3$, we observe from the application of Lemma 4.6 (ii) on $H_{2}=(V_{2},E_{2})$ with $E_{2}=\left\{e_{1,1},e_{1,2}\right\}$ and $V_{2}=\left\{v\in V\mid v\in E_{2}\right\}$ that there exist pairwise distinct $i_{2},i_{3},i_{4}\in\mathbb{N}$ such that $\ell(\left\{v_{1,2},v_{2,2}\right\})=\left\{\left\{i_{2},i_{3}\right\},\left\{i_{3},i_{4}\right\}\right\}$. Moreover, since $\ell(e_{1,1}),\ell(e_{1,2})\in L_{G_{H,\ell}}$ we find by Lemma 3.10 (ii) distinct numbers $i_{1},i_{5}\in\mathbb{N}\setminus\left\{i_{2},i_{3},i_{4}\right\}$ satisfying $\ell(\left\{v_{1,1},v_{2,1}\right\})=\left\{\left\{i_{1},i_{2}\right\},\left\{i_{1},i_{4}\right\}\right\}$ and $\ell(\left\{v_{1,3},v_{2,3}\right\})=\left\{\left\{i_{2},i_{5}\right\},\left\{i_{4},i_{5}\right\}\right\}$. Therefore, conditions (i)(a), (i)(b) and (i)(c) hold for $k=3$. Now, suppose $n\geq 4$ and the statement holds for some particular $3\leq k<n$. Since $\ell(\left\{v_{1,k},v_{2,k}\right\})=\left\{\left\{i_{2},i_{k+2}\right\},\left\{i_{k+2},i_{4}\right\}\right\}$ we can find by Lemma 4.6 (i) a positive integer $i_{k+3}\in\mathbb{N}\setminus\left\{i_{1},\ldots,i_{k+2}\right\}$ satisfying the relation $\ell(\left\{v_{1,{k+1}},v_{2,{k+1}}\right\})=\left\{\left\{i_{2},i_{k+3}\right\},\left\{i_{k+3},i_{4}\right\}\right\}$}. Thus, by our induction hypothesis (i)(a), (i)(b) and (i)(c) hold for $k+1$, which shows the desired statement. As a consequence, we can find for the loop labelings $\ell_{1}$ and $\ell_{2}$ pairwise distinct numbers $i_{1},\ldots,i_{n+2}\in\mathbb{N}$ and $j_{1},\ldots,j_{n+2}\in\mathbb{N}$, respectively, satisfying conditions (i)(a)(i)(c) for $k=n$. Thus, we see that $V_{G_{H,\ell_{1}}}=\left\{i_{1},\ldots,i_{n+2}\right\}$ as well as $V_{G_{H,\ell_{2}}}=\left\{j_{1},\ldots,j_{n+2}\right\}$. Moreover, from the conditions (i)(a)(i)(c) we immediately obtain that the function $f\colon V_{G_{H,\ell_{1}}}\to V_{G_{H,\ell_{2}}}$ defined by $f(i_{k})\coloneqq j_{k}$ for $k=1,\ldots,n+2$ fulfills property (2), which shows that the induced graphs $G_{H,\ell_{1}}$ and $G_{H,\ell_{2}}$ are isomorphic.